OpenStudy (anonymous):

find n if 4P(n,3)=5P(n-1,3)

OpenStudy (aum):

Is this permutations problem? What is the definition of 4P(n,3) ?

OpenStudy (anonymous):

yes, it is.

OpenStudy (anonymous):

i think 4P(n,3) is a set

OpenStudy (callisto):

By definition \[P(n,r)= \frac{n!}{(n-r)!}\] So, for P(n,3), we have \(\frac{n!}{(n-3)!}\), which can be rewritten as \[\frac{n!}{(n-3)!} = \frac{n(n-1)!}{(n-3)!}=\frac{n(n-1)(n-2)!}{(n-3)!}=\frac{n(n-1)(n-2)(n-3)!}{(n-3)!}\]Then we can cancel the one with factorial, and we have \[P(n,3) = n(n-1)(n-2)\]Can you do something similar for P(n-1,3)?

OpenStudy (anonymous):

i'm still kind of lost..sorry

OpenStudy (callisto):

Which part do you not understand?

OpenStudy (anonymous):

i was wondering what happened to the coefficient of "4P" and i can't really figure out how you arrived with n(n-1)!

OpenStudy (callisto):

I think 4P(n,3) = 4 * P(n,3) Recall n! = n * (n-1) * (n-2) * ... * 3 * 2 * 1 So, (n-1)! = (n-1) * (n-2) * (n-3) * ... * 3 * 2 * 1 Combining the two, we have n! = n * (n-1) * (n-2) * (n-3) * ... * 3 * 2 * 1 = n * (n-1)!

OpenStudy (aum):

Example: 7! = 7 x 6! because 7! = 7 x 6 x 5 x 4 x 3 x 2 x 1 6! = 6 x 5 x 4 x 3 x 2 x 1 So 7! = 7 x 6!

OpenStudy (aum):

Is something still not clear?

OpenStudy (anonymous):

i tried solving for (n-1,3), this is what i got \[\frac{ (n-1)(n-1)! }{ (n-3)! } = \frac{ (n-1)(n-1)(n-2)(n-3)! }{ (n-3)! } =(n-1)^2(n-2)\] \[P(n-1,3)=(n-1)^2(n-2)\]

OpenStudy (callisto):

Oh! For \((n-1)!\), we'll have \[(n-1)! \]\[= (n-1) *(n-2)! \]\[= (n-1)*(n-2)*(n-3)! \]\[= (n-1)*(n-2)*(n-3)*(n-4)!\]

OpenStudy (aum):

\[ P(n,r)= \frac{n!}{(n-r)!} \\ P(n-1,r)= \frac{(n-1)!}{(n-1-r)!} \\ P(n-1,3)= \frac{(n-1)!}{(n-4)!} = ~~?\\ \]

OpenStudy (anonymous):

P(n-1,3)=(n-1)(n-2)(n-3) ?

OpenStudy (aum):

correct.

OpenStudy (aum):

You need to solve: 4n(n-1)(n-2) = 5(n-1)(n-2)(n-3)

OpenStudy (aum):

It is a cubic equation and so it will have three values of n that will satisfy the equation.

OpenStudy (anonymous):

can i just divide both sides by (n-1) and (n-2)?

OpenStudy (aum):

Can't do that because we are solving the equation for n and so when n = 1 or n = 2 we will be dividing by zero.

OpenStudy (aum):

Instead you can say when n = 1, the LHS is zero and so is RHS. So n = 1 is one solution. Same thing with n = 2. There is one more n value (other than 1 or 2) that you need to find. Now you can divide both sides by (n-1) and (n-2) because we are looking for an n value other than 1 or 2.

OpenStudy (anonymous):

\[4n^3-12n^2+8n=5n^3-30n^2+40n-30\]

OpenStudy (aum):

Oh, no need to do that! You have common factors on both sides and the solution can be easily found as stated in my previous reply.

OpenStudy (anonymous):

oh,okay. sorry. so, the value of n is 15?

OpenStudy (aum):

n = 1 n = 2 n = 15

OpenStudy (aum):

Sorry, I meant those are solutions to the cubic equation. But you have to check the original equation to see if there are any extraneous solution.

OpenStudy (anonymous):

okay. thank you for your help! :D

OpenStudy (aum):

The original problem doesn't make sense when n = 1 or n = 2 and so they are extraneous solutions. So I think n = 15 is the answer even though Wolfram lists all three answers.